Specially trained dogs are used to detect explosives in the border control. In a random testing of 100 items, trained dogs detected trace of explosives in 4 of them. a) Find the standard deviation and the standard error estimate for the proportion of detected trace for explosives. * ANSWER USING EXCEL FUNCTIONS *   

b) Find the 95% confidence interval for the population proportion of detected trace for explosives. * ANSWER USING EXCEL FUNCTIONS *

c) Find the 90% confidence interval for the population proportion of detected trace for explosives. * ANSWER USING EXCEL FUNCTIONS *

- Thank you in advance
Answer #1
Proportion P =4/100 0.04
Standard deviation =SQRT(0.04*0.96) 0.195959
Standard error =SQRT(0.04*0.96/100) 0.019596
Margin of error =1.96 * SQRT(0.04*0.96/100) 0.038408
95% CI =(P-ME , P+ME) 0.001592 0.078408
Margin of error =1.645 * SQRT(0.04*0.96/100) 0.032235
90% CI = (P-ME , P+ME) 0.007765 0.072235

a.

Standard deviation :

\sqrt{\hat{p} \: (1-\hat{p})} =0.195959

Stnadard error

\sqrt{\frac{\hat{p} \: (1-\hat{p})}{n}} =0.0195959

b.

95% confidence interval is (0.001592, 0.078408)

c.

90% confidence interval is ( 0.007765 , 0.072235 )